OpenStudy (anonymous):

A company sells candy bars. The table below shows the total weight of a box with 1,2,3, and 4 candy bars Number of candy bars in box| Total weight ________________________________________ 1 | 9.5 2 |15.5 3 |21.5 4 |27.5

OpenStudy (anonymous):

@ranga

OpenStudy (jessiegonzales):

So what is the question?

OpenStudy (anonymous):

I am supposed to write an algebraic rule that can be used to determine the total weight of the box for any number of candy bars.

OpenStudy (anonymous):

Well since you add 6 each time you add a candy bar, the algebraic equation would be 3.5 + 6x = Y(Total weight of the box) Where x is the number of candy bars. You get 3.5 by finding the Y-intercept, subtracting 6 from (1, 9.5) to get (0,3.5). The y-intercept symbolizes the weight of the box.

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

How would I find the Y intercept?

OpenStudy (anonymous):

From a graph? and what 2 points would I use?

jimthompson5910 (jim_thompson5910):

3.5 + 6x = y is the same as y = 6x + 3.5 That second equation is of the form y = mx+b slope: m = 6 y-intercept: b = 3.5 which is the point (0,3.5)

jimthompson5910 (jim_thompson5910):

you can use any two points you want to graph the line

OpenStudy (anonymous):

ok, Thanks!

OpenStudy (anonymous):

How did you come up with 3.5?

jimthompson5910 (jim_thompson5910):

do you know how to find the slope m = 6 ?

OpenStudy (anonymous):

Slope is Rise/Run right?

OpenStudy (anonymous):

How do you find the slope?

jimthompson5910 (jim_thompson5910):

using the slope formula m = (y2 - y1)/(x2 - x1)

OpenStudy (anonymous):

So, if I picked (1,3) and (2,3) that wouldn't work...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(1,3) (5,6) Won't work either

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

I can't figure this out @zzr0ck3r I picked 2 random points but they don't work

OpenStudy (zzr0ck3r):

what is your question?

OpenStudy (zzr0ck3r):

when he says pick pionts he does not mean pick random points, he means pick random x values plug that in to get the y value, now you have a point

OpenStudy (anonymous):

but how did he get 3.5?

OpenStudy (zzr0ck3r):

letting x = 0, we see y = 3.5 (0,3.5) is one point letting x = 1 we see y = 9.5 so (1,9.5) is another

OpenStudy (anonymous):

How did he get 3.5 to write this equation y = 6x + 3.5

OpenStudy (zzr0ck3r):

1 -> 9.5 2 -> 15.5 3-> 21.5 so its growing by 6 every time, so to find out what we get when x is zero (i.e. the y intercept) we take away 6 from 9.5 (because its 9.5 at 1)to get 3.5.

OpenStudy (zzr0ck3r):

or take a point (1,9.5) you have y=6x+b plug in your point 9.5=6+b b=3.5

OpenStudy (anonymous):

but to get the point (1,9.5) you needed the equation y = 6x + 3.5

OpenStudy (anonymous):

So I will follow your first suggestion